# Function theory on the annulus in the dp-norm

### v1

@@ Abstract.
where $10<\delta<$ 1. The class of functions in question arises in the early work of R. G. Douglas and V. I. Paulsen on the rational dilation of a Hilbert space operator $T$ to a normal operator with spectrum in $\partial R_{\delta}$. Their work suggested the following norm $\|\cdot\|_{\mathrm{dp}}$ on the space $\mathrm{Hol}(R_{\delta})$ of holomorphic functions on $R_{\delta}$,
@@
Our central result is a generalization of the classical realization formula, for $\varphi\in\mathcal{S}$,
$\varphi(z)=A+Bz(1-Dz)^{-1}C,\ \text{for}\ z\in\mathbb{D},$
to functions from $\mathcal{S}_{\mathrm{dp}}$. A second result is a Pick interpolation theorem for functions in $\mathcal{S}_{\mathrm{dp}}$ that is analogous to Abrahamse’s Interpolation Theorem for bounded holomorphic functions on a multiply-connected domain. For a tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of distinct interpolation nodes in $R_{\delta}$, we introduce a special set $\mathcal{G}_{\mathrm{dp}}(\lambda)$ of positive definite $n\times n$ matrices, which we call DP Szegő kernels. The DP Pick problem $\lambda_{j}\mapsto z_{j},j=1,\dots,n$, is shown to be solvable if and only if,
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0\;\text{ for all}\;g\in\mathcal{G}_{\mathrm{dp}}(\lambda).$
@@ 1. Introduction
$g_{\rho}(\lambda_{i},\lambda_{j})=\sum_{m=-\infty}^{\infty}\frac{(\bar{\lambda}_{i}\lambda_{j})^{m}}{\rho+\delta^{2m}},\quad\text{for}\ 1\leq i,j\leq n.$
@@
$k_{\mathcal{D}}(\lambda,\mu)=-\frac{1}{\bar{\mu}\lambda}\log(1-\bar{\mu}\lambda).$
@@
Hence, we see from equations (1.4) and (1.5) that $\|\varphi\|_{\mathrm{dp}}\leq\left(2+\frac{1+\delta}{1-\delta}\right)\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}$ for $\varphi\in{\rm Hol}(R_{\delta})$. On the other hand, trivially, $\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}\leq\|\varphi\|_{\mathrm{dp}}$, and so the dp and supremum norms on ${\rm Hol}(R_{\delta})$ are equivalent. Thus,
@@
and $STS^{-1}$ dilates to a normal operator with spectrum contained in the boundary $\partial R_{\delta}$. This result is a kind of Nagy dilation theorem for the annulus. In the scalar case a slightly stronger result than the inequality (1.5) had been obtained earlier by A. Shields [23, Proposition 23], with the smaller constant $2+\sqrt{\frac{1+\delta}{1-\delta}}$ on the right hand side. Shields asked whether the constant $2+\sqrt{\frac{1+\delta}{1-\delta}}$ could be replaced by a quantity that remains bounded as $\delta\to 1$. This question was answered in the affirmative by C. Badea, B. Beckermann and M. Crouzeix [7] and subsequently a better constant was established by M. Crouzeix [11].
@@
$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0\ \ \text{ and }\ \ \left[\left(1-\overline{\frac{\delta}{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]\geq 0.$ (1.9)
@@
We observe that $\mathcal{G}_{\mathrm{dp}}(\lambda)$ consists of the gramians $[\langle e_{j},e_{i}\rangle]_{i,j=1}^{n}$ for all bases $e_{1},\dots,e_{n}$ of an $n$-dimensional Hilbert space $\mathcal{H}$ such that the operator $T$ on $\mathcal{H}$ defined by $Te_{j}=\lambda_{j}e_{j}$ for $j=1,\dots,n$ is a Douglas-Paulsen operator. This and related facts are described in Section 4.
@@ 2. The dp and sup norms on ${\rm Hol}(\mathbb{D})$ and ${\rm Hol}(R_{\delta})$
In this section we describe connections between the Banach algebra ${\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ and the Crouzeix conjecture. We will show in Proposition 2.11 that there is a large class of functions $\varphi\in{\rm Hol}(R_{\delta})$, such that
@@
A simple example of a function $f\in{\rm Hol}(R_{\delta})$ for which the last relation fails is $f(z)=z+\frac{\delta}{z}$, which satisfies
@@
for some $\delta$ such that $10\leq\delta<$ 1. Note that any elliptical domain can be identified via an affine self-map of the plane with an elliptical domain of the form $G_{\delta}$ for some $\delta\in[0,1)$.
@@
The B. and F. Delyon family, $\mathcal{F}_{\mathrm{bfd}}(C)$, corresponding to an open bounded convex set $C$ in $\mathbb{C}$ is the class of operators $T$ such that the closure of the numerical range of $T$, $\overline{W(T)}$, is contained in $C$. By [18, Theorem 1.2-1], the spectrum $\sigma(T)$ of an operator $T$ is contained in $\overline{W(T)}$, and so, by the Riesz-Dunford functional calculus, $\varphi(T)$ is defined for all $\varphi\in{\rm Hol}(C)$ and $T\in\mathcal{F}_{\mathrm{bfd}}(C)$. Therefore, we may consider the calcular norm [^3] A calcular norm on a function space is a norm that is defined with the aid of the functional calculus. For more information on such norms the reader may consult [6, Chapter 9].
@@
In [9], Crouzeix proved $K_{\mathrm{bfd}}\leq 12$ and conjectured that $K_{\mathrm{bfd}}=2$. Subsequently Crouzeix and Palencia [10] proved that $K_{\mathrm{bfd}}\leq 1+\sqrt{2}$.
@@
Note that $\varphi(z)=z$ has bfd-norm equal to $2$ and sup norm on $G_{\delta}$ equal to $1+\delta$. Hence the Crouzeix universal constant $K_{\mathrm{bfd}}\geq 2$. ∎
@@
The inequality $K_{\mathrm{bfd}}\geq 2$ was proved by G. Tsikalas [22]. He considered the domain $A_{R}=\{1/R<|z|<R\}$ where $R>1,$ and the functions $g_{n}$ in ${\rm Hol}(A_{R})$ defined by
$g_{n}(z)=\frac{1}{R^{n}}\left(\frac{1}{z^{n}}+z^{n}\right),\ n=1,2,\dots.$
He showed that
$\limsup_{n\to\infty}\frac{\|g_{n}\|_{\mathrm{dp}}}{\|g_{n}\|_{{\rm H}^{\infty}(A_{R})}}\geq 2,$
from which it follows that $K_{\mathrm{bfd}}\geq 2$.
@@
$\leq\sup_{\|X\|\leq 1}\|\varphi(X)\|$ by the definition of $\text{by the definition of }\mathcal{F}_{\mathrm{dp}}(\delta)$
$=\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}.$ by von Neumann’s inequality (2.13)
@@ 3. Models and realizations of holomorphic functions on $R_{\delta}$
In this section we review some known results on the function theory of holomorphic functions in the dp-norm on an annulas. The models and realizations of holomorphic functions $\varphi:R_{\delta}\to\mathbb{C}$ such that $\|\varphi\|_{\mathrm{dp}}\leq 1$ are presented in [6, Theorem 9.46]. The theorem states the following.
@@
$1-\overline{\varphi(w)}\varphi(z)=(1-\bar{w}z)\langle v^{+}(z),v^{+}(w)\rangle_{\mathcal{N}^{+}}+(\bar{w}z-\delta^{2})\langle v^{-}(z),v^{-}(w)\rangle_{\mathcal{N}^{-}}.$
@@
A positive semi-definite function on a set $X$ is a function $A:X\times X\to\mathbb{C}$ such that, for any positive integer $n$ and any points $x_{1},\dots,x_{n}\in X$, the $n\times n$ matrix $[A(x_{j},x_{i})]_{i,j=1}^{n}$ is positive semi-definite.
@@
$1-\overline{\varphi(\mu)}\varphi(\lambda)=(1-\bar{\mu}\lambda)A(\lambda,\mu)+(1-\overline{\frac{\delta}{\mu}}\frac{\delta}{\lambda})B(\lambda,\mu)$ (3.4)
@@
Definition 3.8.
If $\varphi\in{\rm Hol}(R_{\delta})$ we say $\varphi$ is of finite type if $\varphi\in\mathcal{S}_{\mathrm{dp}}$ and, in addition, $\varphi$ can be represented as in equation (3.4) where $A$ and $B$ have finite rank.
@@
$1-\overline{\varphi(\mu)}\varphi(\lambda)=(1-\bar{\mu}\lambda)\langle v^{+}(\lambda),v^{+}(\mu)\rangle_{\mathcal{N}^{+}}+(\bar{\mu}\lambda-\delta^{2})\langle v^{-}(\lambda),v^{-}(\mu)\rangle_{\mathcal{N}^{-}}.$
@@
for all $\lambda\in R_{\delta}$. Since both $\mathcal{X}$ and $\mathcal{Y}$ are subspaces of $\mathbb{C}\oplus\mathcal{N}$, we may extend $L$ (possibly after enlarging the space $\mathcal{N}$) to an isometry $L:\mathcal{N}\to\mathcal{N}$ (see the discussion in [6, Remark 2.31] for this step). Write $L$ as a block operator matrix
@@
where $E(\lambda)$ is given by equation (3.6). Since $\|D\|\leq 1$ and $\|E(\lambda)\|=\max\{|\lambda|,\frac{\delta}{|\lambda|}\}<1$ for all $\lambda\in R_{\delta}$, it follows that $1-DE(\lambda)$ is invertible for $\lambda\in R_{\delta}$, and hence
@@
Conversely, suppose that $a,\beta,\gamma.D$ are such that $L$ given by equation (3.14) is an isometry on $\mathbb{C}\oplus\mathcal{N}$ and that $\varphi$ is the function on $R_{\delta}$ defined by equation (3.11). Since $1-DE(\lambda)$ is invertible for all $\lambda\in R_{\delta}$ we may define a mapping $u:R_{\delta}\to\mathcal{N}$ by equation (3.12). Then the equations (3.15) hold. They may be written in the form
@@
$1-\overline{z_{i}}z_{j}=(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}+(1-\frac{\delta^{2}}{\bar{\lambda}_{i}\lambda_{j}})b_{ij}$
@@ 4. DP-Szegő kernels and normalized DP-Szegő kernels for the tuple $(\lambda_{1},\dots,\lambda_{n})$
In this section we follow Abrahamse’s idea of using families of kernels to solve Pick interpolation problems.
@@
$[(1-\bar{\mu}\lambda)k(\lambda,\mu)]\geq 0\ \ \text{ and }\ \ [(1-\frac{\delta}{\overline{\mu}}\frac{\delta}{\lambda})k(\lambda,\mu)]\geq 0,\qquad\text{for all}\;\lambda,\mu\in R_{\delta}.$ (4.2)
@@
Let $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ be an $n$-tuple of distinct points in $R_{\delta}$. We denote by $\mathcal{F}_{\rm dp}(\delta,\lambda)$ the set of operators $T$ in the Douglas-Paulsen family $\mathcal{F}_{\mathrm{dp}}(\delta)$ corresponding to the annulus $R_{\delta}$ that act on an $n$-dimensional Hilbert space $\mathcal{H}_{T}$ and satisfy
@@
$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0\ \ \text{ and }\ \ \left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]\geq 0.$ (4.6)
@@
$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0\ \ \text{ and }\ \ \left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]\geq 0.$ (4.8)
@@
By assumption $T$ is a Douglas-Paulsen operator with parameter $\delta$ that acts on an $n$-dimensional Hilbert space $\mathcal{H}_{T}$, $T$ has $n$ linearly independent eigenvectors $e_{1},\dots,e_{n}$ corresponding to the eigenvalues $\lambda_{1},\dots,\lambda_{n}$ respectively and $g_{ij}=\langle e_{j},e_{i}\rangle$ for $i,j=1,\dots,n$. By the definition of the Douglas-Paulsen class, $\|T\|\leq 1$ and $\|\delta/T\|\leq 1$, so that, for any vector $x=\sum_{j=1}^{n}x_{j}e_{j}$, we have
00 0 $\leq\|x\|^{2}-\|Tx\|^{2}$
@@
$\big{[}(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}\big{]}_{i,j=1}^{n}\geq 0.$
Likewise, the relation $\|\frac{\delta}{T}x\|\leq\|x\|$ holds for any vector $x=\sum_{j=1}^{n}x_{j}e_{j}\in\mathcal{H}_{T}$. Therefore, we have
00 0 $\leq\|x\|^{2}-\left\|\frac{\delta}{T}x\right\|^{2}$
@@
$=\sum_{j,i=1}^{n}\xi_{j}\lambda_{j}\bar{\xi}_{i}\bar{\lambda}_{i}\langle\,e_{j},e_{i}\,\rangle$
$=\sum_{j,i=1}^{n}\bar{\lambda}_{i}\lambda_{j}\xi_{j}\bar{\xi}_{i}\,g_{ij}.$ (4.12)
@@
$\|x\|^{2}-\|Tx\|^{2}=\sum_{j,i=1}^{n}(1-\bar{\lambda}_{i}\lambda_{j})\,g_{ij}\xi_{j}\bar{\xi}_{i}.$
@@
Let $\lambda_{1},\ldots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$ and $z_{1},\ldots,z_{n}\in\mathbb{C}$. If the DP Pick Problem 3.17 is solvable, then, for any positive definite $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (4.15)
@@
By assumption, the DP Pick Problem 3.17 is solvable, that is, there exist a function $\varphi\in{\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ with $\|\varphi\|_{\mathrm{dp}}\leq 1$ and satisfying
@@
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. We say that a DP-Szegő kernel $[g_{ij}]\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ is normalized if $g_{ii}=1$ for $i=1,\dots,n$.

### v2

@@ Abstract.
where $0<\delta<1$. The class of functions in question arises in the early work of R. G. Douglas and V. I. Paulsen on the rational dilation of a Hilbert space operator $T$ to a normal operator with spectrum in $\partial R_{\delta}$. Their work suggested the following norm $\|\cdot\|_{\mathrm{dp}}$ on the space $\mathrm{Hol}(R_{\delta})$ of holomorphic functions on $R_{\delta}$,
@@
Our central result is a Pick interpolation theorem for functions in $\mathcal{S}_{\mathrm{dp}}$ that is analogous to Abrahamse’s Interpolation Theorem for bounded holomorphic functions on a multiply-connected domain. For a tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of distinct interpolation nodes in $R_{\delta}$, we introduce a special set $\mathcal{G}_{\mathrm{dp}}(\lambda)$ of positive definite $n\times n$ matrices, which we call DP Szegő kernels. The DP Pick problem $\lambda_{j}\mapsto z_{j},j=1,\dots,n$, is shown to be solvable if and only if,
[…]
@@ 1. Introduction
$g_{\rho}(\lambda_{i},\lambda_{j})=\sum_{m=-\infty}^{\infty}\frac{(\overline{\lambda}_{i}\lambda_{j})^{m}}{\rho+\delta^{2m}},\quad\text{for}\ 1\leq i,j\leq n.$
@@
$k_{\mathcal{D}}(\lambda,\mu)=-\frac{1}{\overline{\mu}\lambda}\log(1-\overline{\mu}\lambda).$
@@
Hence, we see from equations (1.4) and (1.5) that
$\|\varphi\|_{\mathrm{dp}}\leq\left(2+\frac{1+\delta}{1-\delta}\right)\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}$
for $\varphi\in{\rm Hol}(R_{\delta})$. On the other hand, see Remark 2.7, $\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}\leq\|\varphi\|_{\mathrm{dp}}$, and so the dp and supremum norms on ${\rm Hol}(R_{\delta})$ are equivalent. Thus,
@@
and $STS^{-1}$ dilates to a normal operator with spectrum contained in the boundary $\partial R_{\delta}$. This result is a kind of Nagy dilation theorem for the annulus. In the scalar case a slightly stronger result than the inequality (1.5) had been obtained earlier by A. Shields [24, Proposition 23], with the smaller constant $2+\sqrt{\frac{1+\delta}{1-\delta}}$ on the right hand side. Shields asked whether the constant $2+\sqrt{\frac{1+\delta}{1-\delta}}$ could be replaced by a quantity that remains bounded as $\delta\to 1$. This question was answered in the affirmative by C. Badea, B. Beckermann and M. Crouzeix [7] and subsequently the better constant $1+\sqrt{2}$ was established by M. Crouzeix and A. Greenbaum [11].
@@
[…]
@@
We observe that $\mathcal{G}_{\mathrm{dp}}(\lambda)$ consists of the gramians $\langle e_{j},e_{i}\rangle]_{i,j=1}^{n}$ for all bases $e_{1},\dots,e_{n}$ of an $n$-dimensional Hilbert space $\mathcal{H}$ such that the operator $T$ on $\mathcal{H}$ defined by $Te_{j}=\lambda_{j}e_{j}$ for $j=1,\dots,n$ is a Douglas-Paulsen operator. This and related facts are described in Section 4.
@@ 2. The dp and sup norms on ${\rm Hol}(\mathbb{D})$ and ${\rm Hol}(R_{\delta})$
In this section we describe connections between the Banach algebra ${\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ and the Crouzeix conjecture. We will prove in Proposition 2.11 that there is a large class of functions $\varphi\in{\rm Hol}(R_{\delta})$, such that
@@
In Example 2.9 below we show that the last relation fails to hold for the function $\varphi\in{\rm Hol}(R_{\delta})$ defined by $\varphi(z)=z+\frac{\delta}{z}$ for $z\in R_{\delta}$. In fact $\varphi$ satisfies
$\|\varphi\|_{\mathrm{dp}}=2\ \text{and}\ \|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}=1+\delta.$
@@
for some $\delta$ such that $0\leq\delta<1$. Note that any elliptical domain can be identified via an affine self-map of the plane with an elliptical domain of the form $G_{\delta}$ for some $\delta\in[0,1)$.
@@
The B. and F. Delyon family, $\mathcal{F}_{\mathrm{bfd}}(C)$, corresponding to an open bounded convex set $C$ in $\mathbb{C}$ is the class of operators $T$ such that the closure of the numerical range of $T$, $\overline{W(T)}$, is contained in $C$. By [19, Theorem 1.2-1], the spectrum $\sigma(T)$ of an operator $T$ is contained in $\overline{W(T)}$, and so, by the Riesz-Dunford functional calculus, $\varphi(T)$ is defined for all $\varphi\in{\rm Hol}(C)$ and $T\in\mathcal{F}_{\mathrm{bfd}}(C)$. Therefore, we may consider the calcular norm [^3] A calcular norm on a function space is a norm that is defined with the aid of the functional calculus. For more information on such norms the reader may consult [6, Chapter 9].
@@
In [9], Crouzeix proved $K_{\mathrm{bfd}}\leq 12$ and conjectured that $K_{\mathrm{bfd}}=2$. Subsequently Crouzeix and Palencia [10] proved that $K_{\mathrm{bfd}}\leq 1+\sqrt{2}$. Still more recently Crouzeix and Kressner [12] showed that $W(T)$ is a complete $(1+\sqrt{2})$-spectral set for $T$.
@@
Note that $\varphi(z)=z$ has bfd-norm equal to 22 and sup norm on $G_{\delta}$ equal to $1+\delta$. Hence the Crouzeix universal constant $K_{\mathrm{bfd}}\geq 2$. ∎
@@
In [23] G. Tsikalas proved a result about the annulus as a $K$-spectral set. We restate his result in the notation of this paper as follows. Let $K(\delta)$ denote the smallest constant such that $R_{\delta}$ is a $K(\delta)$-spectral set for any bounded linear operator $T\in\mathcal{F}_{\rm dp}(\delta)$. He used the functions $g_{n}$ in ${\rm Hol}(R_{\delta})$ defined by
$g_{n}(z)=\frac{\delta^{n}}{z^{n}}+z^{n},\quad\text{for}\ n=1,2,\dots,$
to show that $K(\delta)\geq 2$, for all $\delta\in(0,1)$.
@@
$\leq\sup_{\|X\|\leq 1}\|\varphi(X)\|\$$\text{by the definition of }\mathcal{F}_{\mathrm{dp}}(\delta)$
$=\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}.\$by von Neumann’s inequality (2.13)
@@ 3. Models and realizations of holomorphic functions on $R_{\delta}$
In this section we review some known results on the function theory of holomorphic functions in the dp-norm on an annulus. The models and realizations of holomorphic functions $\varphi:R_{\delta}\to\mathbb{C}$ such that $\|\varphi\|_{\mathrm{dp}}\leq 1$ are presented in [6, Theorem 9.46]. The theorem states the following.
@@
$1-\overline{\varphi(w)}\varphi(z)=(1-\overline{w}z)\langle v^{+}(z),v^{+}(w)\rangle_{\mathcal{N}^{+}}+(\overline{w}z-\delta^{2})\langle v^{-}(z),v^{-}(w)\rangle_{\mathcal{N}^{-}}.$
@@
A positive semi-definite function on a set $X$ is a function $A:X\times X\to\mathbb{C}$ such that, for any positive integer $n$ and any points $x_{1},\dots,x_{n}\in X$, the $n\times n$ matrix $A(x_{j},x_{i})]_{i,j=1}^{n}$ is positive semi-definite.
@@
$1-\overline{\varphi(\mu)}\varphi(\lambda)=(1-\overline{\mu}\lambda)A(\lambda,\mu)+(1-\frac{\delta}{\overline{\mu}}\frac{\delta}{\lambda})B(\lambda,\mu)$ (3.4)
@@
The following theorem is stated as [6, Theorem 9.54]. For the convenience of the reader we shall give a full proof here.
@@
$1-\overline{\varphi(\mu)}\varphi(\lambda)=(1-\overline{\mu}\lambda)\langle v^{+}(\lambda),v^{+}(\mu)\rangle_{\mathcal{N}^{+}}+(\overline{\mu}\lambda-\delta^{2})\langle v^{-}(\lambda),v^{-}(\mu)\rangle_{\mathcal{N}^{-}}.$
@@
for all $\lambda\in R_{\delta}$. Since both $\mathcal{X}$ and $\mathcal{Y}$ are subspaces of $\mathbb{C}\oplus\mathcal{N}$, we may extend $L$ (possibly after enlarging the space $\mathcal{N}$) to a unitary operator $L:\mathcal{N}\to\mathcal{N}$ (see the discussion in [6, Remark 2.31] for this step). Write $L$ as a block operator matrix
@@
where $E(\lambda)$ is given by equation (3.6). Since $\|D\|\leq 1$ and
$\|E(\lambda)\|=\max{\left\{|\lambda|,\frac{\delta}{|\lambda|}\right\}}<1\ \text{ for all}\ \lambda\in R_{\delta},$
it follows that $1-DE(\lambda)$ is invertible for $\lambda\in R_{\delta}$, and hence
@@
Conversely, suppose that $a,\beta,\gamma,D$ are such that $L$ given by equation (3.13) is a unitary operator on $\mathbb{C}\oplus\mathcal{N}$ and that $\varphi$ is the function on $R_{\delta}$ defined by equation (3.10). Since $1-DE(\lambda)$ is invertible for all $\lambda\in R_{\delta}$ we may define a mapping $u:R_{\delta}\to\mathcal{N}$ by equation (3.11). Then the equations (3.14) hold. They may be written in the form
@@
$1-\overline{z_{i}}z_{j}=(1-\overline{\lambda}_{i}\lambda_{j})a_{ij}+(1-\frac{\delta^{2}}{\overline{\lambda}_{i}\lambda_{j}})b_{ij}$
@@ 4. DP-Szegő kernels and normalized DP-Szegő kernels for the tuple $(\lambda_{1},\dots,\lambda_{n})$
In this section we follow Abrahamse’s idea of using families of kernels to solve Pick interpolation problems. To this end we shall introduce several objects that depend on an $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of distinct points in $R_{\delta}$. First we consider the set $\mathcal{F}_{\rm dp}(\delta,\lambda)$ of operators on $n$-dimensional Hilbert space with spectrum $\{\lambda_{1},\dots,\lambda_{n}\}$ which belong to the Douglas-Paulsen family $\mathcal{F}_{\rm dp}(\delta)$. Secondly we define DP Szegő kernels for the $n$-tuple $\lambda$. We establish a close connection between these two objects in Propositions 4.9 and 4.10. Thereby, in Section 5 we shall establish a theorem analogous to Theorem 1.2, Abrahamse’s Theorem.
@@
[…]
@@
Let $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ be an $n$-tuple of distinct points in $R_{\delta}$. We denote by $\mathcal{F}_{\rm dp}(\delta,\lambda)$ the family of operators $T$ in the Douglas-Paulsen family $\mathcal{F}_{\mathrm{dp}}(\delta)$ corresponding to the annulus $R_{\delta}$ that act on an $n$-dimensional Hilbert space $\mathcal{H}_{T}$ and satisfy
@@
[…]
@@
[…]
@@
By assumption $T$ is a Douglas-Paulsen operator with parameter $\delta$ that acts on an $n$-dimensional Hilbert space $\mathcal{H}_{T}$, $T$ has $n$ linearly independent eigenvectors $e_{1},\dots,e_{n}$ corresponding to the eigenvalues $\lambda_{1},\dots,\lambda_{n}$ respectively and $g_{ij}=\langle e_{j},e_{i}\rangle$ for $i,j=1,\dots,n$. By the definition of the Douglas-Paulsen class, $\|T\|\leq 1$ and $\|\delta T^{-1}\|\leq 1$, so that, for any vector $x=\sum_{j=1}^{n}x_{j}e_{j}$, we have
$0$ $\leq\|x\|^{2}-\|Tx\|^{2}$
@@
$\big[(1-\overline{\lambda}_{i}\lambda_{j})g_{ij}\big]_{i,j=1}^{n}\geq 0.\$
Likewise, the relation $\|\delta T^{-1}x\|\leq\|x\|$ holds for any vector $x=\sum_{j=1}^{n}x_{j}e_{j}\in\mathcal{H}_{T}$. Therefore, we have
$0$ $\leq\|x\|^{2}-\left\|\delta T^{-1}x\right\|^{2}$
@@
$=\sum_{j,i=1}^{n}\xi_{j}\lambda_{j}\overline{\xi}_{i}\overline{\lambda}_{i}\langle\,e_{j},e_{i}\,\rangle$
$=\sum_{j,i=1}^{n}\overline{\lambda}_{i}\lambda_{j}\xi_{j}\overline{\xi}_{i}\,g_{ij}.$ (4.12)
@@
$\|x\|^{2}-\|Tx\|^{2}=\sum_{j,i=1}^{n}(1-\overline{\lambda}_{i}\lambda_{j})\,g_{ij}\xi_{j}\overline{\xi}_{i}.$
@@
Let $\lambda_{1},\ldots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$ and $z_{1},\ldots,z_{n}\in\mathbb{C}$. If the DP Pick Problem 3.16 is solvable, then, for any positive definite $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0.$ (4.14)
@@
By assumption, the DP Pick Problem 3.16 is solvable, that is, there exist a function $\varphi\in{\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ with $\|\varphi\|_{\mathrm{dp}}\leq 1$ and satisfying
@@
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. We say that a DP-Szegő kernel $g_{ij}]\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ is normalized if $g_{ii}=1$ for $i=1,\dots,n$.